OpenStudy (konradzuse):

Find the arc length of y = x^(3/2) on [1,4].

OpenStudy (amistre64):

\[\int_C ds\to \int_{1}^{4}(\sqrt{1+(y')^2})\ dx\]

OpenStudy (konradzuse):

cannot really see the integral.. \[\sqrt{1 + y ^{2}} ?\]

OpenStudy (amistre64):

\[\Large \int_C ds\to \int_{1}^{4}(\sqrt{1+(y')^2})\ dx\]

OpenStudy (amistre64):

i can make it in color too :)

OpenStudy (konradzuse):

what colors are we talking about? red? green? blue?

OpenStudy (konradzuse):

so y' = 3x^(2/3)/2?

OpenStudy (amistre64):

\[\huge \color{green}{\int_C ds}\to \int_{1}^{4}(\color{#ff23cc}{ \sqrt{1+(y')^2}})\ dx\]

OpenStudy (amistre64):

3/2 - 2/2 = 1/2

OpenStudy (konradzuse):

\[\int\limits_{1}^{4}\sqrt{1 + 3x^(2/3)/2} dx\]

OpenStudy (amistre64):

not 2/3 ... 1/2; which makes the ^2 nice and purrty

OpenStudy (konradzuse):

oh woops :P soo tired.

OpenStudy (konradzuse):

so then it's going to be 3/2^1/2 but the ^2 negates the 1/2... so it will be sqrt(1-3x/2)dx

OpenStudy (amistre64):

sooo close; (3/2)^2 = 9/4 {S} sqrt(1+ 9x/4) dx

OpenStudy (konradzuse):

ah poo I give up :p

OpenStudy (amistre64):

lol

OpenStudy (konradzuse):

not really :P

OpenStudy (amistre64):

\[\int \sqrt{1+\frac{9}{4}x}\ dx\] \[\int \sqrt{\frac{4+9x}{4}}\ dx\] \[\frac{1}2\int \sqrt{4+9x}\ dx\]

OpenStudy (konradzuse):

tricky tricky

OpenStudy (amistre64):

if we had a 9 wed be set to rock ... \[\frac{1}2\frac 99\int \sqrt{4+9x}\ dx\] \[\frac{1}{18}\int 9(4+9x)^{1/2}\ dx\]

OpenStudy (amistre64):

almost got it nailed down

OpenStudy (amistre64):

that 1/2 comes from a 3/2 so we have to catch it when it pops out with a 2/3

OpenStudy (amistre64):

nah, i missed it; the 2/3 is afterwards

OpenStudy (amistre64):

\[\frac{1}{18} \int9(4+9x)^{1/2}\ dx=\frac{1}{18}\frac{2}{3}\frac{(4+9x)^{3/2}}{3/2}=\frac{4}{162}(4+9x)^{3/2}\] maybe :)

OpenStudy (amistre64):

just a wee bit askewed

OpenStudy (konradzuse):

sorry I went to get a bite to eat, let me review P

OpenStudy (konradzuse):

where did this 9/9 come from? or why exactly?

OpenStudy (amistre64):

when that function inside gets integrated; the chain rules gonna pop out a 9

OpenStudy (amistre64):

since there is no 9 there we need to borrow one from 9/9

OpenStudy (amistre64):

unless you wanna go thru the whole u sub thing where u = 4+9x, du = 9 dx; du/9 = dx \[\frac{1}{2*9}\int (u)^{1/2}\ du\]

OpenStudy (amistre64):

either way, we need to borrow a 9 to get where we want to go with it

OpenStudy (konradzuse):

that's how I learned it, u sub....

OpenStudy (konradzuse):

ooo I is tired.... I got the rest of this. Thanks very much sir/maam have a great night, thanks for all the help tonight

OpenStudy (amistre64):

this one ..

OpenStudy (konradzuse):

damn yuou beat me to it, yues.

OpenStudy (amistre64):

\[\int \sqrt{1+\frac{9}{4}x}\ dx\] is the setup so its just a matter if integrating over the interval

OpenStudy (konradzuse):

is this going to be the same setup for any eq? sqrt(x'^2? + y'^2)?

OpenStudy (amistre64):

yes, you measure the length of any curve in the same way; if its in 3diminsion you add in a z'^2 but yes; distance is distance

OpenStudy (konradzuse):

Ah okay makes sense... I always find my biggest problem with this stuff is trying to understand why I'm doing it. Professors like to give out eq but not explain them, and not explain how a problem can use multiple ones...

OpenStudy (amistre64):

yeah, there is alot of handwaving at times since the proofs behind the intuitions can get complicated

OpenStudy (amistre64):

 its best just to remember this picture and understand that the length of the curve is adding up the total sum of /\s that we measure

OpenStudy (amistre64):

\[\sum \Delta s^2 = \sum (\Delta x^2+\Delta y^2)\] \[\sum \Delta s = \sum \sqrt{\Delta x^2+\Delta y^2}\] as these values go into infinitesimals we get \[\int ds = \int \sqrt{dx^2+dy^2}\]

OpenStudy (konradzuse):

so x' is 1?

OpenStudy (konradzuse):

if y = x^(3/2) you would need to take the ln of both sides and switch around to get x right? Or is there an easier way tro get it?

OpenStudy (amistre64):

\[x'=\frac{dx}{d* }\]where * is the independant variable when \[x'=\frac{dx}{dx}\to x'=1\]

OpenStudy (konradzuse):

okay so it will always be 1 then?

OpenStudy (konradzuse):

if we are doing a dx problem...

OpenStudy (amistre64):

no, for example:\[if\ x'=\frac{dx}{dt}\]then we have no way of determing if x'=1

OpenStudy (amistre64):

IF we are doing a problem where x IS the independant variable; then x'=dx/dx = 1 in this problem; we get dx/dx so yes

OpenStudy (konradzuse):

hmmm?

OpenStudy (amistre64):

spose y was the INDEPENDANT variable in this problem; then: x' = dx/dy and y' = dy/dy=1

OpenStudy (konradzuse):

so if we are given a problem with dx x would be independant, if it's dt t is the inependant, and dty would be y independant?